# How Accurately Do Large Language Models Understand Code?

### v1

@@ Abstract.
This paper presents the first large-scale empirical investigation into the ability of LLMs to understand code. Inspired by mutation testing, we use an LLM’s ability to find faults as a proxy for its deep understanding of code. This approach is based on the insight that a model capable of identifying subtle functional discrepancies must understand the code well. We first inject faults in real-world programs and ask the LLM to localize them, ensuring the specifications are sufficient for fault localization. Next, we apply semantic-preserving code mutations (SPMs) to the faulty programs and test whether the LLMs still locate the faults, verifying their confidence in code understanding. We evaluate nine popular LLMs on $575,000$ debugging tasks automatically sourced across $670$ Java and $637$ Python programs. We find that LLMs lose the ability to debug the same bug in 81% of the faulty programs when SPM are applied, indicating their shallow understanding of code and reliance on code features irrelevant to semantics. We also find the LLMs understand the code earlier in the program more than the one later. This suggests that LLMs’ code comprehension remains tied to lexical and syntactic features due to traditional tokenization designed for natural languages, which overlooks code semantics.
@@ 1. Introduction
Evaluations. We operationalize our evaluation framework to assess the code comprehension abilities of $9$ state-of-the-art LLMs, including open-source, closed-source, reasoning, coding, and general-purpose models. Using four faults and six semantic-preserving mutations of varying strengths, application locations, and sizes, we automatically generate exactly 575,000 debugging tasks sourced from $637$ Python and $670$ Java programs. These tasks span 189 million lines of code (LOC), amounting to approximately 2.9 billion tokens being evaluated. We find that, on average, LLMs fail to identify faults in $74\%$ of debugging tasks in cases where at least one LLM correctly detected the same faults. More interestingly, LLMs fail to recognize the same faults they identified in phase 1, in 81% of the debugging tasks, when semantic-preserving mutations (SPMs) are introduced. For instance, function shuffling mutation reduces the debugging accuracy of LLMs by 83% percent. This provides strong evidence that LLMs’ ability to understand code semantics is overly influenced by non-functional properties of code. LLMs are most sensitive to inducing dead code SPM and least affected by misleading variable name mutations.
Further analysis reveals that the location of faults and the presence of semantic-preserving mutations play a vital role in LLMs’ ability to understand code. LLMs tend to have better comprehension of code located earlier in the program. Specifically, 60% of the faults correctly localized by LLMs are in the first 25% of a program’s lines, while only 13% are detected in the last 25%.
Overall, across $575K$ experiments, we find that LLMs still struggle to recognize fine-grained code semantics, limiting their utility in post-development tasks, which are key components of the software development stack (O’Dell, 2017; Planning, 2002; Britton et al., 2013). Thus, claims such as “LLMs replacing software engineers” are highly exaggerated. Moreover, we believe a significant effort from the software engineering community is necessary to accurately assess LLMs’ capabilities in software engineering tasks, and our study takes a step in this direction.
@@ 5. Results
We analyze the performance of the LLMs across various dimensions, including fault type, fault location, language, SPM type, SPM strength, SPM location, and LLM category. We create a total of 575K debugging tasks for these LLMs, spanning 189 million lines of code, and 2.9 billion tokens. Table 4 documents the experiment statistics.
@@
Total Prompts Generated 893,007
Total Debugging Tasks 575,560
@@
Total LOC Tested 188,670,568 LOC
Total Token Analyzed 2,978,723,984
@@
APIs Gemini, Anthropic, OpenAI
@@ 5.1. RQ1: LLMs’ Code Comprehension
We analyze the accuracy of individual LLMs in finding the faulty line of code. As mentioned below, verifying if the specifications are adequate enough to distinguish the impact of the injection is highly challenging and nearly infeasible at the scale of our experiments. To automate this process, we perform majority voting to ensure the quality of the specification and exclude faulty programs to measure fault localization where all LLM did not correctly localize the faults. Table 3 presents these results. We measure accuracy in these experiments as the ratio of successful debugging tasks to the total number of faulty programs. Overall, fault detection accuracy in Java programs is significantly lower than in Python programs with the exception of Claude. For example, for fault type MisplacedReturn, we find GPT-4o accurately localizing fault in 74.62% of faulty Python programs, whereas this accuracy drops down to only 36.37% for Java programs. Java is statically typed with relatively more semantic information encoded in the syntactic structure of the program than Python, which is dynamically typed. Therefore, LLMs have more opportunities to extract partial semantics from the code. However, the results of all other LLMs except Claude contradict this. Since Claude is the only model to contradict this, it is likely caused by a difference in training datasets for Claude where Claude performs well overall on Java compared to Python. However, Claude performs better in Java than in Python, as expected. Moreover, Claude is the best-performing LLM for Java in Phase I.
@@
For Incorrect Boolean Logic, most LLMs provide higher fault localization accuracy on Java programs. The same behavior is seen with closed-source LLMs (i.e., Gemini and GPT) for fault OffByOne. A common property between the two is that they will likely influence the branching decision. Since Python uses indents to separate code blocks compared to braces in Java, LLMs struggle with the code branching logic in Python. Java explicitly expresses code branching with tokens like "{" and "}" that LLM successfully picks and thus captures the branching logic better. Therefore, faults that directly impact the code branching in Java are localized with higher accuracy by most LLMs. In OffByOne faults, open-source LLMs have lower accuracy when the fault was injected in Java compared to Python, which can be attributed to the generally low reasoning capabilities of LLMs with verbose and imperative languages like Java as compared to the declarative nature of Python.
@@
Across different fault types, the highest detection accuracy achieved by any LLM ranges from $3.25\%$ to $87.78\%$ for Python programs and from $3.33\%$ to $73.26\%$ for Java programs, highlighting a gap of more than $84\%$ in the best detection accuracies. Among all fault types, OperatorSwap and MisplacedReturn are the most easily detected, with accuracies of $87.78\%$ and $83.33\%$ for Python, respectively. This is also observed with the Java dataset for OperatorSwap and MisplacedReturn with accuracies of $73.26\%$ and $70.83\%$ respectively. This is expected because both faults introduced textually noticeable changes in a program path that are most likely to cause an explicit behavioral change (e.g., terminating a method earlier). In contrast, the other two faults introduce changes that implicitly affect the branching logic (e.g., not executing the last iteration of a loop).
@@
We identify the top-performing models in each category. Notably, Gemini 1.5 Pro leads in fault detection performance, demonstrating high code comprehension, followed closely by OpenAI’s GPT‑4o. The highest fault detection accuracies come from closed‑source models, suggesting that proprietary LLMs benefit from extensive training data, deployments on hardware needed to host the full‑scale models, and optimizations. Among open‑source LLMs, Qwen2.5‑coder:7B, Deepseek v3, and Phi‑4:14B perform best, with Qwen2.5‑coder achieving the highest fault detection accuracy (16.75%) for the OffByOne fault and Deepseek v3 achieving the highest fault detection accuracy (16.75%) for IncorrectBooleanLogic in the Python dataset. Deepseek v3, Qwen2.5‑coder:7B, and Phi‑4:14B outperform other open‑source models. Although these models are not specialized for code-specific tasks, they benefit from being trained on a larger and more recent corpus of data compared to older generic models like Llama 3.1. This fresh training data grants it a richer understanding of modern coding practices and language nuances. In contrast, Qwen‑QwQ, despite its impressive 32B parameter size and strong reasoning capabilities, appears to overthink code-related tasks and misanalyze the fault (Shah et al., 2024).
@@ 5.2.1. Fault Localization After Applying Mutations
Figure 6 summarizes the results, averaged across all fault types and SPMs. We measure accuracy in these experiments as the ratio of successful debugging tasks in Phase II to those in Phase I. Almost every model showed a notable drop in fault localization accuracy when SPMs are present. Open-source models such as Llama3.1:8B and Qwen experienced the largest declines, whereas closed-source models like Claude, GPT-4o and Gemini demonstrated greater resilience. This performance gap is largely due to the models’ reliance on surface-level cues such as variable names, comments, and code structure, which they use to infer underlying semantics. When these cues are manipulated without altering the functionality, the models’ learned representations become less reliable. In contrast, closed-source models benefit from extensive fine-tuning on high-quality, diverse datasets, which helps them build deeper semantic understanding and robustness against such perturbations.
@@ 5.2.2. Language-specific Performance under Mutations
For testing with the Deepseek v3 model, we utilized a portion of our originally created dataset. In particular, we use $5\%$ of the debugging tasks from the entire dataset for deepseek due to the relatively slow average response time of 23 seconds per API call. Given this, analyzing the entire dataset would be time-prohibitive. Additionally, since LLMs’ accuracy is low in Phase I, we applied a majority voting approach, including only those debugging tasks detected by more than four LLMs in the $5\%$ sample. Deepseek v3’s results are shown separately in Figure 6, they portray a similar trend that detecting faults in Java is more challenging in comparison to Python and the fault detection accuracy drops further in presence of SPMs.
@@ 5.3.3. Effect of Mutation Strength on LLM’s Code Comprehension
Figure 9 presents the results for each mutation type with varying strength levels (1 to 8), where strength represents the number of times an SPM was applied within a single program. For most LLMs, we find a linear decline in average fault localization accuracy (highlighted by the red line) as the strength of SPM increases. This trend occurs because, as more SPMs are injected, the accumulated syntactic noise progressively obscures the underlying code semantics, making it harder for the models to extract meaningful contextual cues. At lower strength levels, LLMs can often rely on reliable hints from variable names, comments, and code structure to localize faults correctly. For instance, in Java, the accuracy drops on average from $14.79\%$ at strength 1 to $6.71\%$ at strength 8 across all evaluated models, corresponding to an overall decrease of $8.08\%$ over 7 steps (an average decrease of $1.15\%$ per mutation strength increase). Similarly, in Python, the accuracy falls on average from $43.03\%$ at strength 1 to $27.6\%$ at strength 8 across all evaluated models, which represents an overall decline of $15.43\%$ over 7 steps (an average decrease of $2.2\%$ per mutation strength increase). For example, one of the seed Python programs defines a function called make_tile using the Blender Python API to create a 3D model of a tile. In this program, a fault is injected by doing return None early, preventing the execution of the subsequent parts of the function. At a lower SPM strength with a single dead code line, the LLM (Gemini 1.5 Pro) locates the fault. However, when the strength of the dead code SPM was increased to $4$ by introducing multiple dead code statements, it ultimately obscured the injected fault. As a result, the LLM could no longer accurately identify the line with the fault. Similar results can also be observed for the Deepseek v3 model using the same dataset as in Figure LABEL:fig:Llm_type_accuracy_deepseek. Results for the Deepseek model in Figure 10 show a downward trend in average accuracy as mutation strength increases especially for the Python dataset, revealing the compound effect of SPM intensity on fault detection accuracies. This is a prime example of LLMs lacking the depth in semantic understanding required to find faults in the presence of SPMs.
@@ 5.4. RQ4: Effect of Fault Location in the Code
Figure 11 presents a heatmap summarizing fault localization accuracy across different fault types and their positions within the code. Results show that faults in the first quarter of the code ($0-25\%$) are detected with the highest accuracy, suggesting that LLMs may focus more on initial segments and retain clearer context at the beginning of a program. In particular, MisplacedReturn faults are easily detected in this early section. However, detection accuracy declines as faults appear later in the code, with the $75-100\%$ range exhibiting the lowest success rates across all fault types. This decline is most pronounced for Misplaced Return and OffByOne faults, indicating that these error types are not only difficult to detect overall but become even more challenging when positioned later in the program. These findings suggest that LLMs may lose contextual cues or allocate less attention to code segments appearing farther from the start.
@@ 8. Conclusion
While LLMs are increasingly used in software development, their evaluation remains focused on code generation. This paper conducts a large-scale empirical study assessing LLMs’ code comprehension via debugging. We automatically inject faults and semantic-preserving mutations in existing benchmarks to generate a large amount of debugging tasks for LLMs. Experiments on 575K tasks reveal fundamental weaknesses in LLMs, with non-functional code changes reducing debugging accuracy by 81%, highlighting a shallow understanding of code semantics. We identify key code features that challenge LLMs and expose unique weaknesses, guiding research toward more effective LLM use and potential opportunities to improve accuracy.

### v2

@@ Abstract.
This paper presents the first large-scale empirical investigation into the ability of LLMs to understand code. Inspired by mutation testing, we use an LLM’s ability to find faults as a proxy for its deep understanding of code. This approach is based on the insight that a model capable of identifying subtle functional discrepancies must understand the code well. We first inject faults in real-world programs and ask the LLM to localize them, ensuring the specifications are sufficient for fault localization. Next, we apply semantic-preserving code mutations (SPMs) to the faulty programs and test whether the LLMs still locate the faults, verifying their confidence in code understanding. We evaluate nine popular LLMs on $600,010$ debugging tasks automatically sourced across $670$ Java and $637$ Python programs. We find that LLMs lose the ability to debug the same bug in 78% of the faulty programs when SPM are applied, indicating their shallow understanding of code and reliance on code features irrelevant to semantics. We also find the LLMs understand the code earlier in the program more than the one later. This suggests that LLMs’ code comprehension remains tied to lexical and syntactic features due to traditional tokenization designed for natural languages, which overlooks code semantics.
@@ 1. Introduction
Evaluations. We operationalize our evaluation framework to assess the code comprehension abilities of $9$ state-of-the-art LLMs, including open-source, closed-source, reasoning, coding, and general-purpose models. Using four faults and six semantic-preserving mutations of varying strengths, application locations, and sizes, we automatically generate exactly 600,010 debugging tasks sourced from $637$ Python and $670$ Java programs. These tasks span 196 million lines of code (LOC), amounting to approximately 3.1 billion tokens being evaluated. We find that, on average, LLMs fail to identify faults in $74\%$ of debugging tasks in cases where at least one LLM correctly detected the same faults. More interestingly, LLMs fail to recognize the same faults they identified in phase 1, in 78% of the debugging tasks, when semantic-preserving mutations (SPMs) are introduced. For instance, function shuffling mutation reduces the debugging accuracy of LLMs by 83% percent. This provides strong evidence that LLMs’ ability to understand code semantics is overly influenced by non-functional properties of code. LLMs are most sensitive to inducing dead code SPM and least affected by misleading variable name mutations.
Further analysis reveals that the location of faults and the presence of semantic-preserving mutations play a vital role in LLMs’ ability to understand code. LLMs tend to have better comprehension of code located earlier in the program. Specifically, 50% of the faults correctly localized by LLMs are in the first 25% of a program’s lines, while only 18% are detected in the last 25%.
Overall, across $600K$ experiments, we find that LLMs still struggle to recognize fine-grained code semantics, limiting their utility in post-development tasks, which are key components of the software development stack (O’Dell, 2017; Planning, 2002; Britton et al., 2013). Thus, claims such as “LLMs replacing software engineers” are highly exaggerated. Moreover, we believe a significant effort from the software engineering community is necessary to accurately assess LLMs’ capabilities in software engineering tasks, and our study takes a step in this direction.
@@ 5. Results
We analyze the performance of the LLMs across various dimensions, including fault type, fault location, language, SPM type, SPM strength, SPM location, and LLM category. We create a total of 600K debugging tasks for these LLMs, spanning 196 million lines of code, and 3.1 billion tokens. Table 4 documents the experiment statistics.
@@
Total Prompts Generated 930,949
Total Debugging Tasks 600,010
@@
Total LOC Tested 196,685,313 LOC
Total Token Analyzed 3,105,262,000
@@
APIs Gemini, Anthropic, Deepseek, OpenAI
@@ 5.1. RQ1: LLMs’ Code Comprehension
We analyze the accuracy of individual LLMs in finding the faulty line of code. As mentioned below, verifying if the specifications are adequate enough to distinguish the impact of the injection is highly challenging and nearly infeasible at the scale of our experiments. To automate this process, we perform majority voting to ensure the quality of the specification and exclude faulty programs to measure fault localization where all LLM did not correctly localize the faults. Table 3 presents these results. We measure accuracy in these experiments as the ratio of successful debugging tasks to the total number of faulty programs. For testing with the Deepseek v3 model, we utilized a portion of our originally created dataset. In particular, we use $5\%$ of the debugging tasks from the entire dataset for Deepseek due to the relatively slow average response time of 23 seconds per API call. Given this, analyzing the entire dataset would be time-prohibitive. Additionally, since LLMs’ accuracy is low in Phase I, we applied a majority voting approach, including only those debugging tasks detected by more than four LLMs in the $5\%$ sample. Overall, fault detection accuracy in Java programs is significantly lower than in Python programs with the exception of Claude. For example, for fault type MisplacedReturn, we find GPT-4o accurately localizing fault in 74.62% of faulty Python programs, whereas this accuracy drops down to only 36.37% for Java programs. Java is statically typed with relatively more semantic information encoded in the syntactic structure of the program than Python, which is dynamically typed. Therefore, LLMs have more opportunities to extract partial semantics from the code. However, the results of all other LLMs except Claude contradict this. Since Claude is the only model to contradict this, it is likely caused by a difference in training datasets for Claude where Claude performs well overall on Java compared to Python. Claude is the best-performing LLM for Java in Phase I.
@@
For Incorrect Boolean Logic, most LLMs provide higher fault localization accuracy on Java programs. The same behavior is seen with closed-source LLMs (i.e., Gemini, Claude, and GPT) and open-sourced LLMs like Deepseek v3 for fault OffByOne. A common property between the two is that they will likely influence the branching decision. Since Python uses indents to separate code blocks compared to braces in Java, LLMs struggle with the code branching logic in Python. Java explicitly expresses code branching with tokens like "{" and "}" that LLM successfully picks and thus captures the branching logic better. Therefore, faults that directly impact the code branching in Java are localized with higher accuracy by most LLMs. In OffByOne faults, open-source LLMs have lower accuracy when the fault was injected in Java compared to Python, which can be attributed to the generally low reasoning capabilities of LLMs with verbose and imperative languages like Java as compared to the declarative nature of Python.
@@
Across different fault types, the highest detection accuracy achieved by any LLM ranges from $3.25\%$ to $87.78\%$ for Python programs and from $3.33\%$ to $73.26\%$ for Java programs, highlighting a gap of more than $84\%$ in the best detection accuracies. Among all fault types, OperatorSwap and MisplacedReturn are the most easily detected, with accuracies of $87.78\%$ and $83.33\%$ for Python and $73.26\%$ and $70.83\%$ for Java respectively. This is expected because both faults introduced textually noticeable changes in a program path that are most likely to cause an explicit behavioral change (e.g., terminating a method earlier). In contrast, the other two faults introduce changes that implicitly affect the branching logic (e.g., not executing the last iteration of a loop).
@@
We identify the top-performing models in each category. Notably, Claude 3.7 Sonnet and Gemini 1.5 Pro lead in fault detection performance, demonstrating high code comprehension, followed closely by OpenAI’s GPT‑4o. The highest fault detection accuracies come from closed‑source models, suggesting that proprietary LLMs benefit from extensive training data, deployments on hardware needed to host the full‑scale models, and optimizations. Among open‑source LLMs, Qwen2.5‑coder:7B, Deepseek v3, and Phi‑4:14B perform best, with Qwen2.5‑coder achieving the highest fault detection accuracy (16.75%) for the OffByOne fault and Deepseek v3 achieving the highest fault detection accuracy (25.96%) for IncorrectBooleanLogic in the Python dataset. Deepseek v3, Qwen2.5‑coder:7B, and Phi‑4:14B outperform other open‑source models. Although these models are not specialized for code-specific tasks, they benefit from being trained on a larger and more recent corpus of data compared to older generic models like Llama 3.1. This fresh training data grants them a richer understanding of modern coding practices and language nuances. In contrast, Qwen‑QwQ, despite its impressive 32B parameter size and strong reasoning capabilities, appears to overthink code-related tasks and misanalyze the fault (Shah et al., 2024).
@@ 5.2.1. Fault Localization After Applying Mutations
Figure 6 summarizes the results, averaged across all fault types and SPMs. We measure accuracy in these experiments as the ratio of successful debugging tasks in Phase II to those in Phase I. Almost every model showed a notable drop in fault localization accuracy when SPMs were present. Open-source models such as Llama3.1:8B and Qwen models experienced the largest declines, whereas closed-source models like Claude, GPT-4o, and Gemini demonstrated greater resilience. This performance gap is largely due to the models’ reliance on surface-level cues such as variable names, comments, and code structure, which they use to infer underlying semantics. When these cues are manipulated without altering the functionality, the models’ learned representations become less reliable. In contrast, closed-source models benefit from extensive fine-tuning on high-quality, diverse datasets, which helps them build deeper semantic understanding and robustness against such perturbations.
@@ 5.2.2. Language-specific Performance under Mutations
Deepseek v3’s shows a similar trend that detecting faults in Java is more challenging in comparison to Python and the fault detection accuracy drops further in presence of SPMs.
@@ 5.3.3. Effect of Mutation Strength on LLM’s Code Comprehension
Figure 9 presents the results for each mutation type with varying strength levels (1 to 8), where strength represents the number of times an SPM was applied within a single program. For most LLMs, we find a linear decline in average fault localization accuracy (highlighted by the red line) as the strength of SPM increases. This trend occurs because, as more SPMs are injected, the accumulated syntactic noise progressively obscures the underlying code semantics, making it harder for the models to extract meaningful contextual cues. At lower strength levels, LLMs can often rely on reliable hints from variable names, comments, and code structure to localize faults correctly.
For instance, in Java, the accuracy drops on average from $14.79\%$ at strength 1 to $6.71\%$ at strength 8 across all evaluated models, corresponding to an overall decrease of $8.08\%$ over 7 steps (an average decrease of $1.15\%$ per mutation strength increase). Similarly, in Python, the accuracy falls on average from $43.03\%$ at strength 1 to $27.6\%$ at strength 8 across all evaluated models, which represents an overall decline of $15.43\%$ over 7 steps (an average decrease of $2.2\%$ per mutation strength increase). For example, one of the seed Python programs defines a function called make_tile using the Blender Python API to create a 3D model of a tile. In this program, a fault is injected by doing return None early, preventing the execution of the subsequent parts of the function. At a lower SPM strength with a single dead code line, the LLM (Gemini 1.5 Pro) locates the fault. However, when the strength of the dead code SPM was increased to $4$ by introducing multiple dead code statements, it ultimately obscured the injected fault. As a result, the LLM could no longer accurately identify the line with the fault. Similar results can also be observed for the Deepseek v3 model using the same dataset as in Figure 10. This Figure shows a downward trend in average accuracy as mutation strength increases especially for the Python dataset, revealing the compound effect of SPM intensity on fault detection accuracies. This is a prime example of LLMs lacking the depth in semantic understanding required to find faults in the presence of SPMs.
@@ 5.4. RQ4: Effect of Fault Location in the Code
Figure 11 presents a heatmap summarizing fault localization accuracy across different fault types and their positions within the code. Results show that faults in the first quarter of the code ($0-25\%$) are detected with the highest accuracy, suggesting that LLMs may focus more on initial segments and retain clearer context at the beginning of a program. In particular, OperatorSwap faults are easily detected in this early section. However, detection accuracy declines as faults appear later in the code, with the $75-100\%$ range exhibiting the lowest success rates across all fault types. This decline is most pronounced for Misplaced Return faults, indicating that this error type is not only difficult to detect overall but become even more challenging when positioned later in the program. These findings suggest that LLMs may lose contextual cues or allocate less attention to code segments appearing farther from the start.
@@ 8. Conclusion
While LLMs are increasingly used in software development, their evaluation remains focused on code generation. This paper conducts a large-scale empirical study assessing LLMs’ code comprehension via debugging. We automatically inject faults and semantic-preserving mutations in existing benchmarks to generate a large amount of debugging tasks for LLMs. Experiments on 600K tasks reveal fundamental weaknesses in LLMs, with non-functional code changes reducing debugging accuracy by 78%, highlighting a shallow understanding of code semantics. We identify key code features that challenge LLMs and expose unique weaknesses, guiding research toward more effective LLM use and potential opportunities to improve accuracy.